Given that 3 + i is a root of the equation z
3 3 3z
2 2 8z + 30 = 0, find the
remaining roots.
1
Expert's answer
2020-05-20T19:27:52-0400

Given equation is "z^3 - 3z^2 -8z +30=0" .

Also given "z=3+i" is root of given equation, so conjugate of given root is also root of the given equation.

Thus "z=3-i" is root of equation.

So, "z-3-i \ and \ z-3+i" are two factor of given equation

"\implies (z-3-i)(z-3+i) = (z-3)^2 - i^2 = z^2 - 6z +9 +1 = z^2 -6z+10" is factor of given equation.

Now dividing "z^3 - 3z^2 -8z +30=0" by "z^2 -6z+10" , we got quotient is "z+3" .

So, another root of given equation is "z=-3."


Answer: Remaining root of given equation are "z= 3-i, -3."